# Tailoring the Hyperparameters of a Wide-Kernel Convolutional Neural Network to Fit Different Bearing Fault Vibration Datasets

### v1

@@ 3 Method
The second structural component and the majority of the research however focuses on the wide-kernel CNN and aims to investigate how much dataset properties affect which hyperparameter settings are best. We describe the wide-kernel CNN architecture used in our work and its architectural hyperparameters. Training and testing on seven benchmark datasets gives an indication of how differences in the data can have consequences for hyperparameter tuning. We start off with a grid search which is an exhaustive exploration of different combinations of hyperparameters and for each combination we train and test a CNN on benchmark data. We repeat the grid search seven times in order to include all the benchmark datasets mentioned earlier.
In addition to accepting the benchmark datasets as-is, we also want to get into a position to see what happens when we manipulate the dataset properties ourselves. By manipulating the datasets, we gain fine-grained control over factors that might explain why some hyperparameter settings result in high accuracy while others perform badly. In concrete terms, we edit the sampling frequency of the CWRU dataset by resampling it and we also cut out high frequencies by filtering. Both types of data manipulation are described below. Each modified version of the CWRU data becomes the input to a new grid search over important hyperparameters and produces a new grid of test accuracy scores. Collating these grid searches, we can compare how hyperparameters behave in response to different levels of resampling or filtering. At this point, we have several different types of architecture being tested out on different datasets; for clarity, we illustrate the overall flow of the various steps in Figure 1.
Ultimately we wish to extract useful insights into how to tune hyperparameters. There are actually several related analyses based on the outputs of the grid searches, and we go over each one in turn. Correlations are used to see how much hyperparameter settings generalise across datasets. Feature importance is used to see which hyperparameters matter the most. We have two approaches, the first of which is based on Shapley values and the second of which is based on the d-dimensional earth mover’s score. What these terms mean is laid out towards the end of this section.
@@
3.1 Datasets
In this subsection we describe the most obvious input to our experiments: data. We use seven benchmark datasets, described below along with their main properties of the datasets and our pre-processing steps. These are the seven datasets we used earlier in [1].
3.1.1 CWRU Bearing Dataset
The specific choice for this dataset relates to the frequent usage within the fault detection domain  [15, 47, 49, 67, 68, 69], due to its public availability and its structure, which mirrors real-world industrial applications where data is typically not openly accessible. For training purposes, we split the dataset into just 20% train and 80% test data in order to enhance the complexity of the classification task and reduce computation time. For more detailed information, we refer to the original source [66] and the test rig displayed at https://engineering.case.edu/bearingdatacenter/apparatus-and-procedures, along with Figure 2 which shows an example of the vibration signals.
Figure 2: Illustration of vibration signals retrieved from [12], representing two sensors based on the CWRU Bearing dataset. The red box highlights the duration of a single data sample (sequence).
3.1.2 Paderborn Dataset
3.1.3 Gearbox
3.1.4 MFPT Dataset
3.1.5 XJTU
3.1.6 UoC
3.1.7 SEU
3.2 Method Pt. 1 – Experiments on Alternatives to Wide-Kernel CNN Architecture
A wide-kernel CNN architecture takes centre stage in our research however we also want to say that hyperparameters are an important and unavoidable decision for other neural networks too. Therefore, in addition to the wide-kernel CNN which is our focus we also perform limited experiments on two additional architectures. By encompassing more than one architectural style, our work has a broader basis on which more general conclusions can be built.
@@ 3.1.7 SEU
Our goals for the LSTM and transformer architectures are not the same as those we have for the wide-kernel CNN. With the wide-kernel CNN, our intention is to weigh up the importance of different hyperparameters and familiarise ourselves with each hyperparameter in detail. Our results do not generalise to other architectures which have their own distinct sets of hyperparameters. The extra LSTM and transformer architectures are not included to allow for a detailed analysis but simply to provide evidence that hyperparameter tuning has important consequences for a wide range of state-of-the-art neural networks. We want to show that the dangers of ill-hyperparameterisation are not a consequence of choosing the wide-kernel architecture in particular.
3.2.1 Long Short Term Memory (LSTM) Architecture
@@
3.2.2 Transformer Architecture
@@
3.2.3 Hyperparameter Search on Other Architectures
@@
For the LSTM-based neural network, the values we consider for each hyperparameter are shown in Table 1. The middle value for each hyperparameter is based on what it is set to by default in the code published alongside [78]. The other two options simply expand or shrink the amount by a factor of four. From the space of possibilities defined in Table 1, we take 100 random samples.
Table 1: Hyperparameter values considered for the LSTM network.
@@
The transformer hyperparameters are also confined to a limited set of possibilities, which are shown in Table 2. From these we choose a sample of 100 combinations of values. Like with the LSTM, our choice of default options is inspired by the code published by the authors (alongside [50]).
Table 2: Hyperparameter values considered for the transformer network.
@@
3.2.4 Analysis Techniques
Our analysis of the LSTM and transformer architectures asks how much the accuracy of a network changes when its hyperparameters change. The evidence considered is a sample of 100 LSTMs and 100 transformers chosen at random from the range of hyperparameter options, with each network trained and evaluated on seven benchmark datasets. Based on their hyperparameters, some of the 200 networks perform better and some perform worse. To give an overview of the results, we present to the reader a summary of the distribution of accuracy scores. We report the distribution of test accuracy – the spread of how well the networks performed – as follows. First, for the LSTM we state the minimum, 25th percentile, median, 75th percentile and maximum accuracy achieved after training and testing on CWRU data. We then state what values the LSTM reached on the Gearbox data for the minimum, 25th percentile, median, 75th percentile and maximum (also known as Tukey’s five-number summary). The results proceed to give the corresponding values for all the other benchmark datasets and when using the transformer instead. In total, there are 5 values describing how 2 architectures did on 7 different datasets, with the goal being to give an idea of the spread of the data and quickly clarifies how much the architectures are affected by their hyperparameters and by changes in the data. We are also curious how well the default hyperparameter values used in past literature fare. Therefore we also report the accuracy obtained when using the default settings.
3.3 Method Pt. 2 – Focusing on the Wide-Kernel CNN Architecture
@@ 3.3.1 Overview of Architecture
Our architecture contains two convolutional layers, each utilising several combinations of stride, kernel size and filters. After that, a convolutional unit (see Figure 5) is repeated three times, resulting in a total network of five convolutional layers. The three convolutional units all have the same hyperparameter combination in this study, relying on the observation that the deeper layers of the network should have the same properties for optimal learning, as described in previous work [33, 34, 15]. After each convolutional layer, the network utilises local average pooling to decrease the vector size of the convolutional output with length $T$ divided by two, resulting in a pooled output length of $\frac{T}{2}$. At last, the reduced output from the convolutional layers is fed to two fully connected layers and the last one functions as the classification layer. Figure 5 shows the overall network architecture with two exemplary time series as input.
@@ 3.3.2 Hyperparameter Search
Earlier work on the wide-kernel CNN identified seven architectural hyperparameters [2]. These were explored using a grid search strategy in which each hyperparameter is restricted to a limited set of values and every possible way of combining these is tested. For example, every kernel size in the first layer is tried out with every stride in layer 1, and each of these combinations is tried out with every kernel size in layer 2, etc. Since they differ from one another, the set of options tried for any particular hyperparameter is not the same as for other hyperparameters. Table 3 shows the options. Looking at all the mixtures that can be created, there are $5\times 3\times 6\times 2\times 6\times 2\times 6=12960$ combinations included in the grid search.
Previous work also found that three hyperparameters in particular are important: kernel size in the first layer, number of filters in the first layer and number of filters in layers 3-5 (these are indicated in italics in 3) [2, 1]. We use these hyperparameters as a basis for new grid searches. The reason we are performing new grid searches is that we want to find out more about the impact of manipulating the properties of the data by filtering or resampling. Our approach to filtering and resampling the data is described in the next subsections. By manipulating the dataset we create new versions of it. On each version, we perform a grid search considering the values shown in Table 3. The data manipulation strategies and the goals of performing them are explained in more detail in the next subsection.
@@
Table 3: Architectural hyperparameters with the values they take in grid search.
@@
3.3.3 Analysis Techniques on Seven Benchmark Datasets
@@
If possible, we would like to recommend effective default values for the seven architectural hyperparameters of our wide-kernel CNN architecture. Recommendations would be easy to make if we find that a single value performs best across all the seven benchmark datasets. Therefore, we present a simple table that shows the average accuracy obtained on each dataset when using each hyperparameter value. The table compares different options for a hyperparameter, showing if there is a generally best option or if instead there is no obvious default value that performs well on all datasets.
@@
3.3.4 Resampling
@@
To investigate the impact of the sampling rate on hyperparameters such as the kernel size of the first layer, we make multiple copies and resample them by a factor of 2, 4, 8 and 16. Starting off with data sampled at 48 kHz, this results in versions sampled at 24 kHz, 12 kHz, 6 kHz and 3 kHz. On each version of the data, we run a grid search over the hyperparameters which previous work showed were the most important. The grid search results can be analysed to compare what hyperparameter settings work best at different sampling rates.
@@ 3.3.5 Filtering
It should be noted that filtering in this way does not impact the sampling rate or the size of a window of vibration data, it simply removes content which occurs at high frequencies. This makes it easy to compare the results when performing a grid search over the most important hyperparameters. A grid search can be performed on each filtered version of the CWRU dataset, and the results are directly comparable since the same architectures are applied to data which has exactly the same size and shape.
3.3.6 Analysis Techniques on Resampled and Filtered Data
We analyse the resulting information in multiple ways. First, we look at correlations between different versions of the dataset, which gives an indication of the impact that resampling had. If resampling had no impact at all, then we would expect to have a perfect correlation of 1.0, and values lower than this would imply that resampling has an influence on which hyperparameter combinations are stronger and which are weaker.
Afterwards, we look at how important different hyperparameters are. We do this by initially asking a related question of whether the test accuracy of a network can be predicted from the hyperparameters it has. If the accuracy can be predicted, we can hone in on the specifics of how much each hyperparameter contributes to the predictability.
@@ 3.3.6 Analysis Techniques on Resampled and Filtered Data
Correlation analysis: By resampling the CWRU data, we create multiple versions of the dataset. One obvious question that arises is whether a combination of hyperparameters that works well on one version of the data will also work well on another. A similar question could be asked about the worst-performing hyperparameters. In general, this is a question of correlation: to what extent do the accuracy scores obtained on one dataset version correlate to the scores on data resampled to a different rate? We compute and then display the correlations, which gives an insight into how much the hyperparameters are affected by the resampling/filtering. One possibility is that certain hyperparameter combinations are simply better than others, and it does not matter if the properties of the data (such as sampling rate) change. In the extreme, this would result in a correlation of 1.0 between all versions of the dataset. However, if smaller correlations are found, then it would suggest that the properties of the data impact which hyperparameters are best, or that there is some degree of randomness in the accuracy scores that deflates the correlations.
@@
Box plots: One of the key questions in our research is whether the optimal kernel size in the first layer changes when the data is either resampled or filtered. Box plots provide a straightforward view onto how well different kernel sizes tend to work. We plot the distribution of accuracy scores for each kernel size when applied to each version of the data.
Feature importance based on Shapley values: After performing a grid search over different values that the hyperparameters can take, we are left with a set of accuracy scores. Each entry in the results records the specific values that were used for the three hyperparameters manipulated in the grid search, plus the trained network’s accuracy on test data. We use these four pieces of information to investigate the relative importance of each hyperparameter. The analysis proceeds by analysing one grid search at a time, before comparing across grid searches. Remember that a grid search represents a certain level of resampling or filtering, e.g. it tells us what happens when different hyperparameter values are tried out on CWRU data resampled from 48 kHz to 24 kHz. Another grid search covers CWRU data resampled from 48 to 12 kHz, and so on.
‘Importance’, of course, can mean many things. Here, we provide one particular approach to quantifying the importance of hyperparameters. Our approach asks how much each hyperparameter seems to contribute to the final accuracy of a network. If a hyperparameter is important, then setting it correctly tends to result in a noticeably higher accuracy. If a hyperparameter is not important, then even changing it drastically will not cause any increase or decrease in how accurate the network is.
It is difficult to directly state how much a hyperparameter influences the accuracy of the network because this can be contingent based on other hyperparameters or could be subject to a nonlinear relationship which is hard to identify using basic statistical analysis. Our past work showed that nonlinear regression methods were noticeably better at the task of predicting how accurate networks will be. In order to model the obscure, nonlinear relationships between hyperparameters and accuracy, we use a multilayer perceptron. It is a simple neural network – ours contains two hidden layers of 32 neurons – which is told the how the hyperparameters of a wide-kernel CNN have been configured and is trained to predict what accuracy those hyperparameters lead to. The outputs from grid searches provide the data used for training the MLP.
Once the MLP is trained, it contains information about the importance of different hyperparameters. In order to probe the MLP further we train versions which have access to a hyperparameter (i.e., are told what value the hyperparameter has) and versions which do not. If the versions without access to a hyperparameter perform much worse then the hyperparameter is seen as important. Shapley values provide a framework for fairly combining information from all the possible permutations of inclusions/exclusions of hyperparameters [84]. By visualising these Shapley values, we can look into the relative importance of each single hyperparameter, measured in terms of how much it contributes to reducing the MLP regressor’s Mean Absolute Error (MAE). This is the method we employed earlier in [2], although there we applied it to raw benchmark data rather than filtered and resampled CWRU data.
Calculating the Shapley value of a hyperparameter is done by permuting the order of all hyperparameters and then performing two regressions (using the MLP): one with the current hyperparameter included and one without. The difference in error between these two regressions indicates the hyperparameter’s contribution to the overall model’s performance. By averaging over all possible permutations, we obtain the precise Shapley value. For an approximation, a set number of permutations can be chosen at random. For the sake of computational efficiency, we use 100 random permutations. As seen in [85], Shapley values tend to converge quite quickly to a reasonable approximation, and so we do not feel it is necessary to run all permutations. Following the notation in [85], the calculation can be written down follows:
$\varphi_{i}=\frac{1}{\lvert N\rvert!}\sum_{\mathcal{O}\in\pi(\lvert N\rvert)}[v(Pre^{i}(\mathcal{O})\cup\{i\})-v(Pre^{i}(\mathcal{O}))]$ (3)
where $\varphi_{i}$ is the Shapley value of feature $i$, $|N\rvert$ is the number of features, $\mathcal{O}$ is a randomly-chosen permutation of the features, $v(s)$ is the average test performance (mean absolute error or MAE) of a regression using only the set of features $s$, and $Pre^{i}(\mathcal{O})$ is the set of features that appear earlier than feature $i$ in the permutation $\mathcal{O}$.
Feature importance based on d-dimensional earth mover’s distance: We want to understand how much the accuracy of the neural network changes depending on specific settings of the hyperparameters. To do this, we first look at how to summarise a single setting of a single hyperparameter, such as giving the filter in the first layer a width of 256. When the first layer has a filter of width 256, many possibilities remain for the other hyperparameters. Within our grid search, all the possibilities are exhaustively tested, the upshot of which is that we have a collection of many tested networks that all have a first layer filter width of 256. The test accuracy scores from a collection of networks, taken together, form a distribution. Some accuracy scores will be more frequent than others.
The idea of our analysis is to take different distributions and compare them. We can compare the distribution of accuracy scores when the first layer has a filter size of 256 to when the filter size is 128, for example. If two distributions are wildly different, it clearly shows that changing the hyperparameter has caused a change in what accuracy you can expect from the network.
To summarise this information, we use a metric called the d-dimensional Earth Mover’s Distance (EMD) for each hyperparameter [86]. The EMD essentially measures the difference between two probability distributions, reflecting how much one needs to be adjusted to become the other. Higher EMD values indicate greater dissimilarity. The EMD is the same as the Wasserstein difference when comparing two distributions, but it can be extended to compare more than two distributions (hence ‘d-dimensional’). Analysing all the values that a hyperparameter can take, we come up with a single EMD that summarises the hyperparameter overall. With EMD, we can assign a single number to each hyperparameter, representing its impact on accuracy scores. This is a method we used earlier in [1], although there we applied it to raw benchmark data rather than filtered and resampled CWRU data.
It is important to note that some hyperparameters are tested with more values than others are. For instance, we might test five kernel sizes but only three strides and this makes the EMD values somewhat inconsistent and difficult to interpret. To make the results more usable we add an extra processing step that begins by calculating a baseline EMD for each hyperparameter. We do this by calculating the EMD for surrogate data in which the accuracy scores have been shuffled and therefore the relationships between hyperparameters and accuracy have been destroyed. Shuffling multiple times and recalculating the EMD allows us to estimate the average EMD due to chance. When we present the EMD for each hyperparameter we make the numbers more comparable and meaningful by dividing by the baseline.
3.4 Summary of Datasets Used

### v2

@@ III Method
The second structural component of the method and the majority of the research however focuses on the wide-kernel CNN and aims to investigate how much dataset properties affect which hyperparameter settings are best. We describe the wide-kernel CNN architecture used in our work and its architectural hyperparameters. Training and testing on seven benchmark datasets gives an indication of how differences in the data can have consequences for hyperparameter tuning.
In addition to accepting the benchmark datasets as-is, we also want to get into a position to see what happens when we manipulate the dataset properties ourselves. By manipulating the datasets, we gain fine-grained control over factors that might explain why some hyperparameter settings result in high accuracy while others perform badly. In concrete terms, we edit the sampling frequency of the CWRU dataset by resampling it and we also cut out high frequencies by filtering.
At this point, we have several different types of architecture being tested out on different datasets; for clarity, we illustrate the overall flow of the various steps in Figure 1.
@@
III-A Datasets
The most obvious input to our experiments is data. Different types of fault, machinery and recording set-up are possible in fault detection and we therefore expect that datasets can differ from one another. We want to account for variations in the data and so we assembled a selection of seven benchmarks to train and test out CNNs, the same seven that we used earlier in [1]. All of the datasets are benchmarks used for fault detection experiments in industrial machinery, focusing particularly on bearings and gearboxes. All of the datasets are described in more detail in Appendix A.
The CWRU bearing dataset derives from a structured experimental setup which mirrors real-world conditions with various fault locations and severity levels. There are minor imbalances in the class due to differing machine operating conditions. The Gearbox dataset is used to simulate real-time diagnostics, with balanced data for classifying healthy versus faulty gearboxes under varying load conditions. The MFPT dataset explores faults in machinery with diverse fault types and unbalanced data due to different sampling frequencies. The XJTU datasets provide extensive run-to-failure data for bearing faults makes it suitable for analysing the remaining useful life. The UoC dataset focuses on gear faults with a balanced setup, targeting multi-class classification tasks through single-channel vibrational data. Lastly, the SEU datasets offer insights into bearings and gearboxes with comprehensive multivariate data across various fault states and operational conditions.
Overall, these datasets are instrumental for developing and evaluating fault detection algorithms in machinery, each contributing unique features to test.
III-B Method Pt. 1 – Experiments on Alternatives to Wide-Kernel CNN Architecture
A wide-kernel CNN architecture takes centre stage in part 2 of our method, however we also want to say that hyperparameters are an important and unavoidable decision for other neural networks too. Therefore, in addition to the wide-kernel CNN which is our focus we also perform limited experiments on two additional architectures. By encompassing more than one architectural style, our work has a broader basis on which more general conclusions can be built.
@@ III-A Datasets
In this way, the goal of part 1 of our method is to provide evidence that hyperparameter tuning has important consequences for a wide range of state-of-the-art neural networks. We want to show that the dangers of ill-hyperparameterisation are not a consequence of choosing the wide-kernel architecture in particular.
III-B1 Long Short Term Memory (LSTM) Architecture
@@
Like [54], we place an LSTM in an overall architecture that has two CNN components feeding into the LSTM itself. The two CNNs are used in order to achieve a “multi-scale” effect, with one CNN containing smaller kernels and the other containing larger kernels. Both CNNs summarise the raw inputs into a more concise sequence, which is needed because the LSTM layer would be quickly overwhelmed otherwise. The LSTM component is a pair of repeated layers. After the LSTM, fully connected layers lead to a final prediction of what type of fault is present in the machinery. The architecture is visualised in Appendix B-A.
III-B2 Transformer Architecture
@@
The network we use receives incoming signals in a CNN component which leads to a compact sequence that then passes through to the transformer component. The CNN has only two layers and it shrinks the raw signal to a sequence of roughly 1/16th the length. The transformer section of the architecture consists of 4 layers by default. Global average pooling layer is used to consolidate the outputs from the transformer into a single vector. Finally there is a single fully-connected layer which outputs the final prediction of the network. The architecture is visualised in Appendix B-B.
III-B3 Hyperparameter Search on LSTM and Transformer Architectures
@@
For the LSTM-based neural network, the values we consider for each hyperparameter are shown in Table I. The middle value for each hyperparameter is based on what it is set to by default in the code published alongside [54]. The other two options simply expand or shrink the amount by a factor of four. From the space of possibilities defined in Table I, we take 100 random samples.
TABLE I: Hyperparameter values considered for the LSTM network.
@@
The transformer hyperparameters are also confined to a limited set of possibilities, which are shown in Table II. From these we choose a sample of 100 combinations of values. Like with the LSTM, our choice of default options is inspired by the code published by the authors (alongside [38]).
TABLE II: Hyperparameter values considered for the transformer network.
@@
III-B4 Analysis Techniques
Our analysis of the LSTM and transformer architectures asks how much the accuracy of a network changes when its hyperparameters change. The evidence considered is a sample of 100 LSTMs and 100 transformers chosen at random from the range of hyperparameter options, with each network trained and evaluated on seven benchmark datasets. Based on their hyperparameters, some of the 200 networks perform better and some perform worse. When we report the results later, we encapsulate the distribution of accuracy scores (per benchmark) through Tukey’s five-number summary. We are also curious how well the default hyperparameter values used in past literature fare. Therefore we also report the accuracy obtained when using the default settings.
III-C Method Pt. 2 – Focusing on the Wide-Kernel CNN Architecture
@@ III-C1 Overview of Architecture
Our architecture contains two convolutional layers, each utilising several combinations of stride, kernel size and filters. After that, a convolutional unit (see Figure 10) is repeated three times, resulting in a total network of five convolutional layers, as described in previous work [36, 37, 35]. After each convolutional layer, the network utilises local average pooling to decrease the vector size of the convolutional output. At last, the reduced output from the convolutional layers is fed into fully connected layers that classify the vibrations. See Figure 10 in Appendix B for a visual depiction of the architecture, and additional supplementary details.
@@ III-C2 Hyperparameter Search
Earlier work on the wide-kernel CNN identified seven architectural hyperparameters [2]. These were explored using a grid search strategy in which each hyperparameter is restricted to a limited set of values and every possible way of combining these is tested. For example, every kernel size in the first layer is tried out with every stride in layer 1, and each of these combinations is tried out with every kernel size in layer 2, etc. Since they differ from one another, the set of options tried for any particular hyperparameter is not the same as for other hyperparameters. Table III shows the options. Looking at all the mixtures that can be created, there are $5\times 3\times 6\times 2\times 6\times 2\times 6=12960$ combinations included in the grid search.
Previous work also found that three hyperparameters in particular are important: kernel size in the first layer, number of filters in the first layer and number of filters in layers 3-5 (these are indicated in italics in Table III) [2, 1]. We use these hyperparameters as a basis for new grid searches. The reason for new grid searches is that we want to find out more about the impact of manipulating the properties of the data by filtering or resampling. In the next subsections we describe our approach to filtering and resampling the data to create new versions of it On each version, we perform a grid search considering the values shown in Table III.
@@
TABLE III: Architectural hyperparameters with the values they take in grid search.
@@
III-C3 Analysis Techniques on Seven Benchmark Datasets
@@
If possible, we would like to recommend effective default values for the seven architectural hyperparameters of our wide-kernel CNN architecture. Recommendations would be easy to make if we find that a single value performs best across all the seven benchmark datasets. Therefore, we present a simple table that shows the average accuracy obtained on each dataset when using each hyperparameter value. The table compares different options for a hyperparameter, showing if there is a generally best option or if instead there is no value that performs well on all datasets.
@@
III-C4 Multiple Defaults
Similar to previous work [8], we pursue the idea of “multiple defaults”. Rather than providing a single combination of hyperparameters we aim to provide a series of hyperparameter combinations to try out. Our goal is that, after trying $n$ elements of the series, we have as high a chance as possible that at least 1 of the hyperparameter combinations performs well on our data. Different defaults might perform well on different datasets. We now define the task of finding multiple defaults.
For algorithm $A$, let $\Theta$ be a hyperparameter space defined by hyperparameters $\Theta_{1},\Theta_{2},\ldots,\Theta_{n}$, and $\theta$ denote a ‘combination’ that specifies a value for each hyperparameter. Each possible $\theta$ provides all the hyperparameters needed to run the algorithm and can be thought of as a point within $\Theta$.
We intend to find multiple $\theta$ which will be the multiple defaults. To distinguish them, we index them $\theta_{1},\theta_{2},\ldots$ where each $\theta$ is a different combination of hyperparameter settings, i.e., a different point in the hyperparameter space. Each combination is itself composed of multiple values, one value for each hyperparameter. To distinguish the values that compose a single combination, we introduce the following notation. Let hyperparameters within combination $\theta_{x}$ be written as $\theta_{x,1},\theta_{x,2},\ldots,\theta_{x,n}$. For the $x$th hyperparameter combination, the value of the $y$th hyperparameter would be $\theta_{x,y}$.
Our search for multiple defaults will be guided by how well the hyperparameter combinations perform on different benchmark datasets. Let us say there are $l$ different benchmarks. In our research there are seven benchmark datasets, i.e., $l=7$. Each benchmark acts as a function on hyperparameter combinations, written below as: $b_{1}(\theta),b_{2}(\theta),\ldots,b_{l}(\theta)$. Each function takes a hyperparameter combination as an input and returns a real number which is the percentile of how well that combination performed compared to other combinations.
Next we describe the search for multiple defaults. The search proceeds iteratively through multiple steps. At each step of the search, the goal is to identify which hyperparameter combination should be added to the series in order to increase (as much as possible) the maximum performance of any hyperparameter combination seen so far. The maximum performance of hyperparameter combinations $\theta_{1},\theta_{2},\ldots,\theta_{m}$ on benchmark 1 is $max(b_{1}(\theta_{1}),b_{1}(\theta_{2}),\ldots,b_{1}(\theta_{m}))$. At the end of step $m$, we have chosen $m$ hyperparameter combinations. With combinations $1,\ldots,m$ and benchmarks $1,\ldots,l$, the expected best performance of the $m$ combinations is
$\footnotesize\begin{split}E_{m}=\mathbb{E}\Big{|}\max\Bigl{(}b_{1}(\theta_{1}),\ldots,b_{1}(\theta_{m})\Bigr{)},\max\Bigl{(}b_{2}(\theta_{1}),\ldots,b_{2}(\theta_{m})\Bigr{)},\ldots,\\ \max\Bigl{(}b_{l}(\theta_{1}),\ldots,b_{l}(\theta_{m})\Bigr{)}\Big{|}\end{split}$ (2)
The choice of the $m$th hyperparameter combination is the one which increases the expected best performance of the previous $m-1$ combinations the most.
Previously, a greedy algorithm was suggested for finding multiple defaults [8]. The greedy approach is relatively efficient to compute but it does not guarantee the global optimum. By contrast, since we have a completed grid search to work with, we are able to take an exhaustive approach and thus find the globally best-performing series. Because the completed grid search already contains the performance scores (for each combination of hyperparameters on each benchmark) it is possible to execute a brute force algorithm. The computationally expensive step of training and evaluating the CNNs has already been done, making the brute force search a nested loop over the table of results.
In theory, we could keep on adding more and more hyperparameter combinations into our series of multiple defaults. However, in practice we find that the search will identify a relatively small series of hyperparameter combinations that cannot be improved further. Therefore, we terminate our search after $m+1$ steps when we find that $E_{m+1}=E_{m}$, and we return the series found at the end of the first $m$ steps (we return $\theta_{1},\theta_{2},\ldots,\theta_{m}$).
There is one final follow-on analysis which we include in order to evaluate how well the method generalises to new datasets. In this extra analysis we perform multiple runs of the search using leave-one-out validation. When there are $l$ benchmarks, $l-1$ benchmarks are used to determine defaults and then the last benchmark is employed to evaluate their effectiveness. This cycle is repeated $l$ times, ensuring each benchmark is omitted once. Within each repetition we calculate the best performance when applying defaults to the unseen benchmark. Averaging across the $l$ repetitions of the cycle gives an estimated performance on new data.
In summary, our search for multiple defaults leads to $m$ different combinations of hyperparameters, arranged in a sequence. On new data, these combinations can be tried out in turn. If we do not try out all $m$ combinations but only $k$ of them, we still want to maximise the chance of finding a combination that works well. In our research we use seven benchmark datasets to evaluate how well different hyperparameter settings perform. After trying $k$ combinations, for each benchmark we see what did best. The best score per benchmark can be averaged, giving an idea of how quickly the procedure finds at least one successful hyperparameter combination on new data.
III-C5 Resampling
@@
To investigate the impact of the sampling rate on hyperparameters such as the kernel size of the first layer, we make multiple copies of the CWRU data and resample them by a factor of 2, 4, 8 and 16. Starting off with data sampled at 48 kHz, this results in versions sampled at 24 kHz, 12 kHz, 6 kHz and 3 kHz. As resampling increases, the inputs to the neural network get smaller. So we can resample more times before the inputs are too small for our CNN to handle, we start with windows of size 4096 for the resampling experiments, rather than 2048 (the default window size, used elsewhere in this paper, and described in Appendix A). A larger window size results in fewer examples to train from, which makes training more difficult and leads to lower average accuracy. On each version of the data, we run a grid search over the hyperparameters which previous work found most important. The grid search results can be analysed to compare what hyperparameter settings work best at different sampling rates.
@@ III-C6 Filtering
It should be noted that filtering in this way does not impact the sampling rate or the size of a window of vibration data, it simply removes content which occurs at high frequencies. This makes it easy to compare the results when performing a grid search over the most important hyperparameters. A full grid search over hyperparameters can be performed on each filtered version of the CWRU dataset, and the results are directly comparable since the same architectures are applied to data which has exactly the same size and shape.
III-C7 Analysis Techniques on Resampled and Filtered Data
We analyse the resulting information in multiple ways. First, we look at correlations between different versions of the dataset. Afterwards, we look at how well different kernel widths in the first layer perform.
Correlation analysis:
@@ III-C7 Analysis Techniques on Resampled and Filtered Data
By resampling the CWRU data, we create multiple versions of the dataset. One obvious question that arises is whether a combination of hyperparameters that works well on one version of the data will also work well on another. A similar question could be asked about the worst-performing hyperparameters. In general, this is a question of correlation: to what extent do the accuracy scores obtained on one dataset version correlate to the scores on data resampled to a different rate? If resampling had no impact at all, then we would expect to have a perfect correlation of 1.0, and values lower than this would imply that resampling has an influence on which hyperparameter combinations are best. We compute and then display the correlations in a correlation plot, which gives an insight into how much the hyperparameters are affected by the resampling/filtering.
Box plots:
@@
One of the key questions in our research is whether the optimal kernel size in the first layer changes when the data is either resampled or filtered. Box plots provide a straightforward view onto how well different kernel sizes tend to work. We plot the distribution of accuracy scores for each kernel size when applied to each version of the data. This lets us question whether a wide kernel is desirable because of its ability to handle high-frequency noise, or whether another explanation is needed.
III-D Summary of Datasets Used